OpenStudy (anonymous):

if there are n teams in a league and each team plays each other twice in a season, the total number of games is given by the polynomial n^2-n=N. A women's basketball league plays a total of 56 games. How many teams are in the league?

OpenStudy (anonymous):

So let's say t is the number of teams \[t^2-t=56\] Can you solve that?

OpenStudy (anonymous):

56^2?

OpenStudy (ybarrap):

Use the quadratic formula

OpenStudy (ybarrap):

\(t^2-t-56=0\) \(t=\cfrac{-(-1)\pm\sqrt{-1^2-4\times(-56)}}{2}\)

OpenStudy (ybarrap):

http://www.wolframalpha.com/input/?i=t^2-t-56%3D0

OpenStudy (ybarrap):

Make sense?